OpenStudy (anonymous):

If Tucson's average rainfall is 12 1/4 and Yumas is 2 3/5 inches, how much more rain, on an average does Tucson get?

OpenStudy (anonymous):

You can think 12 1/4 is equal to 48/4+1/4=49/4 and 2 3/5 is 10/5+3/5=13/5 So Tucson gets 49/4-13/5= (49*5)/4*5-(13*4)/5*4=245/20-52/20=(245-52)/20=193/20=9,65 more inches of rain than Yuma does, on average.